OpenStudy (anonymous):

Will Fan+Medal :) How many solutions does this system of equations have? y = –3x + 7 y = –3x – 6

OpenStudy (anonymous):

infinitely many 2 1 0

OpenStudy (anonymous):

@pooja195 ? :)

OpenStudy (anonymous):

@sammixboo ? :)

OpenStudy (anonymous):

@DanJS ?

OpenStudy (danjs):

both have the same slope but different y - intercepts

OpenStudy (danjs):

what can you say about that

OpenStudy (anonymous):

That there are more than one?

OpenStudy (danjs):



OpenStudy (danjs):

both have the same slope, but different Y intercepts

OpenStudy (danjs):

The solutions are where the lines intersect, the points (x,y)

OpenStudy (anonymous):

Oh... so would it be two, ore infinitely many?

OpenStudy (danjs):

If two lines have the same slope, and different y intercepts, they are parallel

OpenStudy (danjs):

parallel lines do not intersect. The solution to the system of equations, if any, are where the two lines intersect.

OpenStudy (danjs):

infinitely many solutions describes a system where both equations are the same line. they overlap one another.

OpenStudy (danjs):

parallel lines , are a system of equations that do not intersect, and have no solutions

OpenStudy (anonymous):

Oh :) so the correct answer would be D?

OpenStudy (danjs):

y = -3x+7 y = -3x - 6 trying to solve these, since they are already each solved for y, you can set Y=Y or -3x + 7 = -3x -6 adding 3x to both sides leaves 7 = -6 if you get nonsense like that false statement, then the lines are parallel, and have zero solutions

OpenStudy (danjs):

An example of infinite solutions would be something like y = -6x + 14 y = -6x + 14 obviously the same line, when you try to solve y = y -6x + 14 = -6x + 14 add 6x on both sides 14 = 14 if you get a true statement like that, then the lines are the same line, and overlap, and have infinite solutions

OpenStudy (anonymous):

Thank you so much... you are so helpful :)

OpenStudy (danjs):

welcome, anytime

OpenStudy (danjs):

your medal, just pushed me to 97, woohoo

OpenStudy (anonymous):

Yay :)